If a and b are roots of 2tan^2c+35tanc+2=0 then show that a+b=90°

Grade:11

1 Answers

Utsav Basu
6 years ago
a and b are roots of 2tan2 c + 35tan c + 2 = 0
So,
     tan a * tan b = 2/2 = 1
or, tan a = cot b
or, a + b = 90°